Suppose two dice was rolled. Let Y be the random variable representing the number of even number that occur. Find the values of the random variable Y.

1
Expert's answer
2022-01-24T17:15:50-0500

"P(Y=0)=\frac{9}{36}=\frac{1}{4}."

"P(Y=1)=\frac{18}{36}=\frac{1}{2}."

"P(Y=2)=\frac{9}{36}=\frac{1}{4}."